can anyone help my with this what is proper operaton for this
2(x-3w) = -3(x-w)

x-a=x+a+4+4b can you help?

Answer by sukusu(21)

the proper operation is distributive property
2x-6w= -3x- 3w
see if u can take it from there
x-a=x+a+4+4b
-x
-a=a+4+4b
see if u can take it from there